# 11footnotetext: Equal Contributions

### v4

@@ 4.4 Key Findings
Finding 1: Compliant Residual Policy is able to improve base policy by a large margin. As shown in Fig. 6, [Compliant Residual] policy trained with [On-Policy Delta] data improves the base policy success rate by 60% and 50% on the two tasks respectively. It effectively learns useful corrective strategies from the limited demonstrations. For example, in the book flipping task, the policy learns to pitch the fingers down more before finger insertion to increase the insertion success; in the belt assembly task, the policy learns to correct the height of the belt when misaligning to the pulley slot. Results are best viewed in our supplementary video.
@@
Finding 3: Smooth On-Policy Delta data enables stable residual policy. [Compliant Residual] policy has 45% higher success rate when trained on [On-Policy Delta] data instead of [Take-over-Correction] data on the book flipping task. Residual policy trained with [Take-over-Correction] data sometimes exhibits large noisy motions that trigger task failures, such as retracting the fingers too early in the book flipping task. On the contrary, residual policy trained with [On-policy Delta] data have much smoother action trajectories and better reflect human’s correction intentions, providing suitable magnitudes of corrections.
@@ 4.5 Ablations
We study two important design decisions with ablation studies on the book flipping task.
Training frequency and batch size. One important parameter in DAgger is the batch size between policy updates. With a smaller batch size, the policy is updated more frequently, then new correction data can better reflect the most recent policy behavior. However, DAgger with small batch sizes is known to suffer from catastrophic forgetting [kirkpatrick2017overcoming, goodfellow2013empirical] since it finetunes neural networks on data with non-stationary distribution. Common solutions include retraining the residual policy at the end of DAgger using all available correction data collected from all the intermediate residual policies [wu2025robocopilot]. Another way is to rely on the base policy training data as a normalizer [he2025uncertainty]. In this work, we empirically found that larger batch sizes can effectively stabilize residual training. With batch size = 50, the book flipping task reach 100% with one batch, while the belt assembly task performs better with two batches. We compare our method with a smaller batch size on the book flipping task, where we warm up the residual with 20 episodes of initial correction data, then update every ten more episodes for three times.
@@
5 Conclusion and Discussion
@@
Limitations and Future Work.
The base policy should have a reasonable success rate for the residual policy to learn effectively. From our experiments, we recommend starting to collect correction data for the residual policy when the base policy has at least $10\%\sim 20\%$ success rate. A future direction is to derive theoretical guidelines for the trade-off between the base and residual improvements.
Throughout this work, we use a MLP as the action head of our Compliant Residual Policy and directly regress the actions. Although it works well in our tasks, it may experience difficulty for tasks that involve more distinctive action multi-modalities. More expressive policy formulations, such as Flow Matching [lipman2022flow, black2024pi_0] might be useful for these tasks.
Our data collection system is based on kinesthetic teaching. Although it provides richer data with higher quality than teleoperation as explained in the paper, it may require more labor during training data collection since the demonstrator needs to grasp the handle on the robot.
6 Acknowledgment
We would like to thank Chendong Xin for his support on experiments and rebuttal. We would like to also thank Eric Cousineau, Huy Ha, and Benjamin Burchfiel for thoughtful discussions on the proposed method, thank Mandi Zhao, Maximillian Du, Calvin Luo, Mengda Xu, and all REALab members for their suggestions on the experiment setup and the manuscript. This work was supported in part by the NSF Award #2143601, #2037101, and #2132519, Sloan Fellowship, Toyota Research Institute, and Samsung. We would like to thank Google and TRI for the UR5 robot hardware. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.
@@ Appendix A Technical Appendices and Supplementary Material
A.1 Control Implementation Details
A.1.1 Compliance Control
Compliance is a motor property that describes how motion responds to force. For example, a balloon has high compliance, meaning that they deform significantly under external forces. Industrial robots typically have very low compliance, meaning that their motion tracking has little deviation under external forces.
Compliance control refers to the technique that makes a rigid robot behave softly using feedback control. It is a standard technique widely adopted in the manufacturing industry. It lets robots interact with the environment safely without creating huge forces. With compliance, the robot retreats when it experiences external forces, larger force leads to bigger position deviations. When there is no external force, compliance control tracks the target position accurately just like position-based control. Compliance control allows specification of the exact desired compliance profile, typically described by parameters such as stiffness, damping and inertia.
Impedance control and admittance control are two methods to implement compliance control. The users are free to choose the controller that works with their robot. Robots with good position control accuracy (e.g., most industrial robots) can use admittance control. Robots with low gear ratio (e.g. "quasi-direct-drive" robots) can use impedance control. We use admittance control in this work and open-sourced our implementation at https://github.com/yifan-hou/force_control. A good reference for different compliance control schemes is the “Force Control” chapter of the Handbook of Robotics [villani2016force].
A.1.2 Control Architecture
@@ A.2 Stage-Wise Success Rate
We report the success rate of the book flipping task into three key stages.
@@ A.3 Correction Data Decomposition
As mentioned in the “Training strategy” part of § 3.2, we used two strategies to ensure the residual policy behaves stably around low correction data regions. The first strategy is to include the no correction portion of online data for training and label them with zero residual actions. The second strategy is to collect a few trajectories (15 out of the 50 total correction episodes) in which the demonstrator marks the whole trajectory as correction, even when the correction is small or zero. In practice, we find that the first strategy works better when the base policy is more stable and has a higher success rate, while the second strategy works better otherwise. In our experiments, we use the first strategy for the book flipping task and use the second strategy for the belt assembly task.
A.4 Experiments Compute Resources
@@
A.5 Hardware Design
Our kinesthetic correction hardware setup features a tool interface that allows task-specific tool swapping. For the book flipping task, we designed a customized fork-shaped tool that can easily insert under the books and flip them. For the belt assembly task, we used a standard WSG-50 gripper and fin-ray fingers [chi2024universal]. An interesting future direction is to leverage generative models for automatic manipulator design [xu2024dynamics]. Future work can also incorporate other types of force or tactile sensors, such as capacitive F/T sensors [choi2025coinft] and vision-based tactile sensors [yuan2017gelsight, liu2023enhancing, bauza2024simple].
A.6 Broader Impact
CR-DAgger  contributes to the field of robotics by improving pretrained real-world manipulation policies with a small amount of human correction data. The proposed Compliant Intervention Interface provides an intuitive and safe way for humans to directly interact with robots and correct the robot policy on the spot. We demonstrate significant policy improvements on two real-world contact-rich manipulation tasks, book flipping and belt assembly, which can lead to useful applications in industry and households.

### v5

@@ 4.4 Key Findings from Comparisons
Finding 1: Compliant Residual Policy is able to improve base policy by a large margin. As shown in Fig. 9, [Compliant Residual] policy trained with [On-Policy Delta] data improves the base policy success rate by 60% and 50% on the two tasks respectively. It effectively learns useful corrective strategies from the limited demonstrations. For example, in the book flipping task, the policy learns to pitch the fingers down more before finger insertion to increase the insertion success; in the belt assembly task, the policy learns to correct the height of the belt when misaligned to the pulley slot. Results are best viewed in our supplementary video.
@@
Figure 10: [On-Policy Delta] enables smoother trajectories. Here compares velocity magnitude within $1.5\text{\,}\mathrm{s}$ of the corrections starts/ends. [On-Policy Delta] velocity magnitudes are smaller and more consistent, [Take-Over Correction] has notably larger magnitude and variations, demonstrating that [On-Policy Delta] encourages smoother trajectories.
Figure 11: [On-Policy Delta] introduces less distribution shift. Here we compare the distribution of fingertip trajectories among 1) base policy training data, 2) correction episodes from [On-Policy Delta], and 3) correction episodes from [Take-Over-Correction]. For intuitive visualization, all trajectories are stretched to the same duration and normalized to (0,1). The range of the vertical axis is 0.55m for all three axes. [On-Policy Delta] data’s distribution is better aligned with base policy training data’s distribution than [Take-Over-Correction] data.
Finding 3: Smooth On-Policy Delta data enables stable residual policy. [Compliant Residual] policy has 45% higher success rate when trained on [On-Policy Delta] data instead of [Take-over-Correction] data on the book flipping task. Residual policy trained with [Take-over-Correction] data sometimes exhibits large noisy motions that trigger task failures, such as retracting the fingers too early in the book flipping task. On the contrary, the residual policy trained with [On-policy Delta] data has much smoother action trajectories and better reflects human’s correction intentions, providing suitable magnitudes of corrections. To see the improvements more clearly, we plot robot velocity data around correction starts and ends from 100 episodes of data collections in Fig. 10. [Take-over-Correction] data contains larger robot velocity magnitude right after control authority switches, especially when robot takes over control from human. Moreover, the demonstrator can directly perceive the base policy’s intention and the extent of correction being applied through the resistance force when employing the [On-Policy Delta] kinesthetic teaching style correction, thus preventing correction demonstrations from deviating too much from the base policy’s state-action distribution. Fig. 11 shows that [On-Policy Delta] introduces less distribution shift compared to [Take-over-Correction].
@@ 4.5 Ablations
We study the effect of important design decisions in our method with ablation studies.
Training frequency and batch size. One important parameter in DAgger is the batch size between policy updates. With a smaller batch size, the policy is updated more frequently, then new correction data can better reflect the most recent policy behavior. However, DAgger with small batch sizes is known to suffer from catastrophic forgetting kirkpatrick2017overcoming; goodfellow2013empirical since it finetunes neural networks on data with non-stationary distribution. Common solutions include retraining the residual policy at the end of DAgger using all available correction data collected from all the intermediate residual policies wu2025robocopilot. Another way is to rely on the base policy training data as a normalizer he2025uncertainty. In this work, we empirically found that larger batch sizes can effectively stabilize residual training. With batch size = 50, the book flipping task reach 100% with one batch, while the belt assembly task performs better with two batches. We compare our method with a smaller batch size on the book flipping task, where we warm up the residual with 20 episodes of initial correction data, then update every ten more episodes for three times. We also experiment with multiple batches on the belt assembly task to validate if the improvement is monotonic.
@@
Figure 13: Multi-Batch Update Result. The compliant residual policy trained on two batches of correction data achieves the best result on the belt assembly task, indicating that more batches of correction data do not guarantee better performance.
Finding: More batches does not always bring better performance. Fig. 13 shows the change of success rate on the belt assembly task as the number of batches increases from one to three. The policy trained with three batches of data performs slightly worse than the one trained with two batches, indicating that more correction data does not always bring better performance. This could be caused by distribution shift as we observed different failure modes during each new batch of data collection, and the 50 additional episodes of data was not sufficient to cover the new distribution.
@@
Intention Misinterpretation and tracking error. As explained in Sec. 3, the correction data can misinterpret the human correction behavior when tracking error is large. During development of our method, we observed this phenomenon on the gear insertion task, where the human correction motion is small in magnitude and mostly aligns with the base policy motion, both of which moves towards the pole. As a result, the residual policy learns to pull the gear away from the pole instead of moving it towards the pole, causing a persistent failure. With our fixes that reduces tracking error (equation 2 and 3), the correction data can correctly reflect the human intention and train a residual policy that improves success rate.
The problem does not affect the book flipping task and the belt assembly task even before our fixes are applied. For the book flipping task, the correction motion magnitude is typically much larger than the robot tracking error. The belt assembly task uses minute correction motion around the second pulley, however, the human correction motion is mostly vertical, in which direction the robot has little motion and tracking error.
5 Limitation and Future Work
Base/residual trade-off. The base policy should have a reasonable success rate for the residual policy to learn effectively. From our experiments, we recommend starting to collect correction data for the residual policy when the base policy has at least $10\%\sim 20\%$ success rate. A practical question is whether to spend resources on improving the base policy or the residual policy, given a challenging new task. A future direction is to derive theoretical guidelines for such trade-off between the base and residual improvements.
[…]
More flexible/intuitive data collection interface. Our data collection system is based on kinesthetic teaching. Although it provides richer data with higher quality than teleoperation as explained in the paper, it may require more labor during training data collection since the demonstrator needs to grasp the handle on the robot. Two interesting future work directions include 1) develop teleoperation systems with the same intuitive and smooth correction data collection, and 2) develop intuitive interface for manipulation with more DoFs, such as bi-manual or dexterous hands manipulation.
6 Conclusion
@@
{acks}
We would like to thank Eric Cousineau, Huy Ha, and Benjamin Burchfiel for thoughtful discussions on the proposed method, thank Mandi Zhao, Maximillian Du, Calvin Luo, Mengda Xu, and all REALab members for their suggestions on the experiment setup and the manuscript.
{funding}
This work was supported in part by the NSF Award #2143601, #2037101, and #2132519, Sloan Fellowship, Toyota Research Institute, and Samsung. We would like to thank Google and TRI for the UR5 robot hardware. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.
@@ Appendix A Technical Appendices and Supplementary Material
A.1 Controller Architecture
@@
We choose admittance control to implement compliance in this work, since the Universal Robot we use has high position-tracking accuracy and is not back-drivable. We open-sourced our implementation at https://github.com/yifan-hou/force_control. A more through review of compliance controller implementations can be found in lynch2017modern and villani2016force.
@@ A.2 Stage-Wise Success Rate
We report the success rate of the book flipping task into three key stages. Fig. 14 below is a more detailed version of Fig. 9, which reports all the task success rates by stages.
@@ A.3 Correction Data Decomposition
As mentioned in the “Training strategy” part of § 3.2, we used two strategies to ensure the residual policy behaves stably around low correction data regions. The first strategy is to include the no correction portion of online data for training and label them with zero residual actions. The second strategy is to collect a few trajectories (15 out of the 50 total correction episodes) in which the demonstrator marks the whole trajectory as correction, even when the correction is small or zero. In practice, we find that the first strategy works better when the base policy is more stable and has a higher success rate, while the second strategy works better otherwise. In our experiments, we use the first strategy for the book flipping task and use the second strategy for the other three tasks.
A.4 Comparison of Policy Structures
To further explore alternative residual policy formulations beyond the standard MLP action head used in our experiments, we tested two additional popular policy structures on the belt assembly task: a probabilistic MLP duan2016benchmarking, haarnoja2018soft and a Flow-Matching Transformer lipman2022flow, black2024pi_0. The probabilistic MLP extends the original deterministic MLP by outputting both the mean and log-variance of a Gaussian distribution over actions and sampling actions from this distribution. The Flow Matching Transformer uses a transformer-based conditional flow model that predicts action velocities along a continuous time trajectory, enabling ODE-based sampling over the residual action distribution instead of one-step prediction.
We trained residual policies with these two models using the same data under identical configurations as the original MLP structure on the belt assembly task. The probabilistic MLP achieves a success rate of 100%, and the Flow Matching Transformer achieves 90% (each ablation was evaluated over 10 rollouts). Compared to the success rate of 96.8% achieved by the original MLP, the improvements from more expressive policy structures are inconclusive. The failure modes in this task are relatively homogeneous. We hypothesis that the benefits of expressive policy structure could be more evident for tasks that exhibit more distinctive action multi-modality and leave the more detailed exploration to future work.
A.5 Experiments Compute Resources
@@
A.6 Hardware Design
Our kinesthetic correction hardware setup features a tool interface that allows task-specific tool swapping. For the book flipping task, we designed a customized fork-shaped tool that can easily insert under the books and flip them. For the belt assembly task, we used a standard WSG-50 gripper and fin-ray fingers chi2024universal. An interesting future direction is to leverage generative models for automatic manipulator design xu2024dynamics. Future work can also incorporate other types of force or tactile sensors, such as capacitive F/T sensors choi2025coinft and vision-based tactile sensors yuan2017gelsight, liu2023enhancing, bauza2024simple.